OpenStudy (anonymous):

Rationalize the denominator: (6)/5√3

OpenStudy (anonymous):

multiply it by\[(\frac{ 5\sqrt{3} }{ 5\sqrt{3} })\]

OpenStudy (anonymous):

(30√3)/5√3 ?

OpenStudy (anonymous):